ABSTRACT. In an inner product space X, a cone or a linear variety that is generated by a finite number of linear functionals in the dual space X^{∗} is a Chebyshev set (i) if and only if each of the generating functionals attains its norm, and (ii) if and only if each of them has a “Riesz representer” in X.